a rhombus with a 24 cm diagonal has an area of 120 cm sqaured . find the length of a side of the rhombus
Answer by khwang(438)

Since the area rhombus = product of the two diagonals
Now, 120 = 24 * another diagonal, so another diagonal = 5.

Let x be the length of a side of the rhombus,
we have x^2 = (24/2)^2 + (5/2)^2 = 144 + 25/4 = 601 /4
So, x = sqrt(601)/2